### v3

# Quantum Curriculum Learning

Abstract
Quantum machine learning (QML) requires significant quantum resources to address practical real-world problems. When the underlying quantum information exhibits hierarchical structures in the data, limitations persist in training complexity and generalization. Research should prioritize both the efficient design of quantum architectures and the development of learning strategies to optimize resource usage. We propose a framework called quantum curriculum learning (Q-CurL) for quantum data, where the curriculum introduces simpler tasks or data to the learning model before progressing to more challenging ones. Q-CurL exhibits robustness to noise and data limitations, which is particularly relevant for current and near-term noisy intermediate-scale quantum devices. We achieve this through a curriculum design based on quantum data density ratios and a dynamic learning schedule that prioritizes the most informative quantum data. Empirical evidence shows that Q-CurL significantly enhances training convergence and generalization for unitary learning and improves the robustness of quantum phase recognition tasks. Q-CurL is effective with broad physical learning applications in condensed matter physics and quantum chemistry.
pacs:
Valid PACS appear here
Introduction.— In the emerging field of quantum computing (QC), there is potential to use large-scale quantum computers to solve certain machine learning (ML) problems far more efficiently than classical methods. This synergy between ML and QC has given rise to quantum machine learning (QML) [1, 2], although its practical applications remain uncertain. Classical ML traditionally focuses on extracting and replicating features based on data statistics, while QML is hoped to detect correlations in classical data or generate patterns that are challenging for classical algorithms to achieve [3, 4, 5, 6, 7]. However, it remains unclear whether analyzing classical data fundamentally requires quantum effects. Furthermore, there is a question as to whether speed is the only metric by which QML algorithms should be judged [8]. This suggests a fundamental shift: it is preferable to use QML on data that is already quantum in nature [9, 10, 11, 12, 13, 14].
Figure 1: Overview of two principal methodologies in quantum curriculum learning: (a) task-based and (b) data-based approaches. In the task-based approach, a model ${\mathcal{M}}$, designated for a main task that may be challenging or constrained by data accessibility, benefits from pre-training on an auxiliary task. This auxiliary task is either relatively simpler (left panel of (a)) or has a richer dataset (right panel of (a)). In the data-based approach, we implement a dynamic learning schedule to modulate data weights, thereby emphasizing the significance of quantum data in optimizing the loss function to reduce the generalization error.
The learning process in QML involves extensive exploration within the domain landscape of a loss function. This function measures the discrepancy between the quantum model’s predictions and the actual values, aiming to locate its minimum. However, the optimization often encounters pitfalls such as getting trapped in local minima [15, 16] or barren plateau regions [17]. These scenarios require substantial quantum resources to navigate the loss landscape successfully. Additionally, improving accuracies necessitates evaluating numerous model configurations, especially against extensive datasets. Given the limitation of quantum resources in designing QML models, we must focus not only on their architectural aspects but also on efficient learning strategies.
The perspective of quantum resources refocuses our attention on the concept of learning. In ML, learning refers to the process through which a computer system enhances its performance on a specific task over time by acquiring and integrating knowledge or patterns from data. We can improve current QML algorithms by making this process more efficient. For example, curriculum learning [18], inspired by human learning, builds on the idea of introducing simpler concepts before progressing to complex ones, forming a strategy—a curriculum—that presents easier samples or tasks first. Although curriculum learning has been extensively applied in classical ML [19, 20, 21], its exploration in the QML field, especially regarding quantum data, is still in the early stages. Existing research has primarily examined model transfer learning in hybrid classical-quantum networks [22], where a pre-trained classical model is enhanced by adding a variational quantum circuit. However, there is still limited evidence showing that curriculum learning can effectively improve QML by scheduling tasks and samples.
We explore the potential of curriculum learning using quantum data. We implement a quantum curriculum learning (Q-CurL) framework in two common scenarios. First, a main quantum task, which may be challenging due to the high-dimensional nature of the parameter space or the limitation of data availability, can be facilitated through the hierarchical parameter adjustment of auxiliary tasks. These auxiliary tasks are comparatively easier or more data-rich. However, it is necessary to establish the criteria that make an auxiliary task beneficial for a main task. Second, QML often involves noisy inputs that exhibit a hierarchical arrangement of entanglement or noisy labels, reflecting levels of importance during the optimization process. Recognizing these levels is essential for ensuring the robustness and reliability of QML methods in practical scenarios.
We propose two principal approaches to address the outlined scenarios: task-based Q-CurL [Fig. 1(a)] for the first and data-based Q-CurL [Fig. 1(b)] for the second scenario. In task-based Q-CurL, the curriculum order is defined by the fidelity-based kernel density ratio between quantum datasets. This enables efficient auxiliary task selection without solving each one, reducing data demands for the main task and decreasing training epochs, even if total data requirements stay constant. In data-based Q-CurL, we employ a dynamic learning schedule that adjusts data weights to prioritize quantum data in optimization. This adaptive cost function is broadly applicable to any cost function without requiring additional quantum resources. Empirical evidence shows that task-based Q-CurL enhances training convergence and generalization when learning complex unitary dynamics. Additionally, data-based Q-CurL increases robustness, particularly in noisy-label scenarios, by preventing complete memorization of the training data. This avoids overfitting and improves generalization in the quantum phase detection task. These results suggest that Q-CurL could be broadly effective for physical learning applications.
Task-based Q-CurL.— We formulate a framework for task-based Q-CurL. The target of learning is to find a function (or hypothesis) $h:{\mathcal{X}}\to{\mathcal{Y}}$ within a hypothesis set ${\mathcal{H}}$ that approximates the true function $f$ mapping ${\bm{x}}\in{\mathcal{X}}$ to ${\bm{y}}=f({\bm{x}})\in{\mathcal{Y}}$. To evaluate the correctness of $h$ given the data $({\bm{x}},{\bm{y}})$, the loss function $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is used to measure the approximation error $\ell(h({\bm{x}}),{\bm{y}})$ between the prediction $h({\bm{x}})$ and the target ${\bm{y}}$. We aim to find $h\in{\mathcal{H}}$ that minimizes the expected risk over the distribution $P({\mathcal{X}},{\mathcal{Y}})$:
$R(h):={\mathbb{E}}_{({\bm{x}},{\bm{y}})\sim P({\mathcal{X}},{\mathcal{Y}})}\left[\ell(h({\bm{x}}),{\bm{y}})\right].$ (1)
In practice, since the data generation distribution $P({\mathcal{X}},{\mathcal{Y}})$ is unknown, we use the observed dataset ${\mathcal{D}}={({\bm{x}}_{i},{\bm{y}}_{i})}_{i=1}^{N}\subset{\mathcal{X}}\times{\mathcal{Y}}$ to minimize the empirical risk, defined as the average loss over the training data:
$\hat{R}(h)=\frac{1}{N}\sum_{i=1}^{N}\ell(h({\bm{x}}_{i}),{\bm{y}}_{i}).$ (2)
Given a main task ${\mathcal{T}}_{M}$, the goal of task-based Q-CurL is to design a curriculum for solving auxiliary tasks to enhance performance compared to solving the main task alone. We consider ${\mathcal{T}}_{1},\ldots,{\mathcal{T}}_{M-1}$ as the set of auxiliary tasks. The training dataset for task ${\mathcal{T}}_{m}$ is ${\mathcal{D}}_{m}\subset{\mathcal{X}}^{(m)}\times{\mathcal{Y}}^{(m)}$ ($m=1,\ldots,M$), containing $N_{m}$ data pairs. We focus on supervised learning tasks with input quantum data ${\bm{x}}^{(m)}_{i}$ in the input space ${\mathcal{X}}^{(m)}$ and corresponding target quantum data ${\bm{y}}^{(m)}_{i}$ in the output space ${\mathcal{Y}}^{(m)}$ for $i=1,\ldots,N_{m}$. The training data $\left({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i}\right)$ for task ${\mathcal{T}}_{m}$ are drawn from the probability distribution $P^{(m)}({\mathcal{X}}^{(m)},{\mathcal{Y}}^{(m)})$ with the density $p^{(m)}({\mathcal{X}}^{(m)},{\mathcal{Y}}^{(m)})$. We assume that all tasks share the same data spaces ${\mathcal{X}}^{(m)}\equiv{\mathcal{X}}$ and ${\mathcal{Y}}^{(m)}\equiv{\mathcal{Y}}$, as well as the same hypothesis $h$ and loss function $\ell$ for all $m$.
Depending on the problem, we can decide the curriculum weight $c_{M,m}$, where a larger $c_{M,m}$ indicates a greater benefit of solving ${\mathcal{T}}_{m}$ for improving the performance on ${\mathcal{T}}_{M}$. We evaluate the contribution of solving task ${\mathcal{T}}_{i}$ to the main task ${\mathcal{T}}_{M}$ by transforming the expected risk of training ${\mathcal{T}}_{M}$ as follows:
$R_{T_{M}}(h)$ $={\mathbb{E}}_{({\bm{x}},{\bm{y}})\sim P^{(M)}}\left[\ell(h({\bm{x}}),{\bm{y}})\right]$
$={\mathbb{E}}_{({\bm{x}},{\bm{y}})\sim P^{(m)}}\left[\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}\ell(h({\bm{x}}),{\bm{y}})\right].$ (3)
The curriculum weight $c_{M,m}$ can be determined using the density ratio $r({\bm{x}},{\bm{y}})=\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}$ without requiring the density estimation of $p^{(M)}({\bm{x}},{\bm{y}})$ and $p^{(m)}({\bm{x}},{\bm{y}})$. The key idea is to estimate $r({\bm{x}},{\bm{y}})$ using a linear model $\hat{r}({\bm{x}},{\bm{y}}):={\bm{\alpha}}^{\top}{\bm{\phi}}({\bm{x}},{\bm{y}})=\sum_{i=1}^{N_{M}}\alpha_{i}\phi_{i}({\bm{x}},{\bm{y}}),$ where the vector of basis functions is ${\bm{\phi}}({\bm{x}},{\bm{y}})=(\phi_{1}({\bm{x}},{\bm{y}}),\ldots,\phi_{N_{M}}({\bm{x}},{\bm{y}}))$, and the parameter vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{N_{M}})^{\top}$ is learned from data [23].
[…]
$R_{T_{M}}(h)\approx\dfrac{1}{N_{m}}\sum_{i=1}^{N_{m}}\hat{r}_{\bm{\alpha}}({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})\ell(h({\bm{x}}^{(m)}_{i}),{\bm{y}}^{(m)}_{i}).$ (4)
The parameter vector ${\bm{\alpha}}$ is estimated via the problem of minimizing $\dfrac{1}{2}{\bm{\alpha}}^{\top}{\bm{H}}{\bm{\alpha}}-{\bm{h}}^{\top}{\bm{\alpha}}+\dfrac{\lambda}{2}{\bm{\alpha}}^{\top}{\bm{\alpha}},$ where we consider the regularization coefficient $\lambda$ for $L_{2}$-norm of ${\bm{\alpha}}$. Here, ${\bm{H}}$ is the $N_{M}\times N_{M}$ matrix with elements $H_{ll^{\prime}}=\dfrac{1}{N_{m}}\sum_{i=1}^{N_{m}}\phi_{l}({\bm{x}}_{i}^{(m)},{\bm{y}}_{i}^{(m)})\phi_{l^{\prime}}({\bm{x}}_{i}^{(m)},{\bm{y}}_{i}^{(m)})$, and ${\bm{h}}$ is the $N_{M}$-dimensional vector with elements $h_{l}=\frac{1}{N_{M}}\sum_{i=1}^{N_{M}}\phi_{l}({\bm{x}}_{i}^{(M)},{\bm{y}}_{i}^{(M)})$.
We consider each $\hat{r}({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})$ as the contribution of the data $({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i})$ from the auxiliary task ${\mathcal{T}}_{m}$ to the main task ${\mathcal{T}}_{M}$. We define the curriculum weight $c_{M,m}$ as (see [23] for more details):
$c_{M,m}=\dfrac{1}{N_{m}}\sum_{i=1}^{N_{m}}\hat{r}_{\bm{\alpha}}({\bm{x}}^{(m)}_{i},{\bm{y}}^{(m)}_{i}).$ (5)
We consider the unitary learning task to verify the curriculum criteria based on $c_{M,m}$. We aim to optimize the parameters ${\bm{\theta}}$ of a $Q$-qubit circuit $U({\bm{\theta}})$, such that, for the optimized parameters ${\bm{\theta}}_{\textrm{opt}}$, $U({\bm{\theta}}_{\textrm{opt}})$ can approximate an unknown $Q$-qubit unitary $V$ ($U,V\in\mathcal{U}(\mathbb{C}^{2^{Q}})$).
[…]
$C_{{\mathcal{D}}_{{\mathcal{Q}}}(N)}({\bm{\theta}}):=1-\dfrac{1}{N}\sum_{j=1}^{N}|\braket{\psi_{j}}{V^{\dagger}U({\bm{\theta}})}{\psi_{j}}|^{2}.$ (6)
The parameterized ansatz $U({\bm{\theta}})$ can be modeled as $U({\bm{\theta}})=\prod_{l=1}^{L}U^{(l)}({\bm{\theta}}_{l})$, consisting of $L$ repeating layers of unitaries. Each layer $U^{(l)}({\bm{\theta}}_{l})=\prod_{k=1}^{K}\exp{(-i\theta_{lk}H_{k})}$ is composed of $K$ unitaries, where $H_{k}$ are Hermitian operators, ${\bm{\theta}}_{l}$ is a $K$-dimensional vector, and ${\bm{\theta}}=\{{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{L}\}$ is the $LK$-dimensional parameter vector.
We present a benchmark of Q-CurL for learning the approximation of the unitary dynamics of the spin-1/2 XY model with the Hamiltonian $H_{XY}=\sum_{j=1}^{Q}\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+h_{j}\sigma_{j}^{z}\right)$, where $h_{j}\in{\mathbb{R}}$ and $\sigma_{j}^{x},\sigma_{j}^{y},\sigma_{j}^{z}$ are the Pauli operators acting on qubit $j$. This model is important in the study of quantum many-body physics, as it provides insights into quantum phase transitions and the behavior of correlated quantum systems.
To create the main task ${\mathcal{T}}_{M}$ and auxiliary tasks, we represent the time evolution of $H_{XY}$ via the ansatz $V_{XY}$, which is similar to the Trotterized version of $\exp(-i\tau H_{XY})$ [12]. The target unitary for the main task, $V^{(M)}_{XY}=\prod_{l=1}^{L_{M}}V^{(l)}(\bm{\beta}_{l})\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}}$, consists of $L_{M}=20$ repeating layers, where each layer $V^{(l)}(\bm{\beta}_{l})$ includes parameterized z-rotations RZ (with assigned parameter $\bm{\beta}_{l}$) and non-parameterized nearest-neighbor $\sqrt{i\textup{SWAP}}=\exp(\frac{i\pi}{8}(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}))$ gates. Additionally, we include the fixed-depth unitary $\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}}$ with $L_{F}=20$ layers at the end of the circuit $V^{(l)}$ to increase expressivity. Similarity, keeping the same ${\bm{\beta}}_{l}$, we create the target unitary for the auxiliary tasks ${\mathcal{T}}_{m}$ as $V^{(m)}_{XY}=\prod_{l=1}^{L_{m}}V^{(l)}({\bm{\beta}}_{l})\prod_{l=1}^{L_{F}}V^{(l)}_{\textrm{fixed}}$, with $L_{m}=1,2,\ldots,19$.
Figure 2(a) depicts the average HS distance over 100 trials of ${\bm{\beta}}_{l}$ and $V^{(l)}_{\textrm{fixed}}$ between the target unitary of each auxiliary task ${\mathcal{T}}_{m}$ (with $L_{m}$ layers) and the main task ${\mathcal{T}}_{M}$. We also plot the curriculum weight $c_{M,m}$ in Fig. 2(a) calculated in Eq. (5). Here, we consider the unitary $V_{XY}$ learning with $Q=4$ qubits via the hardware efficient ansatz $U_{\textrm{HEA}}({\bm{\theta}})$  [24, 23] and use $N=20$ Haar random states for input data ${\bm{x}}_{i}^{(m)}$ in each task ${\mathcal{T}}_{m}$. As depicted in Fig. 2(a), $c_{M,m}$ can capture the similarity between two tasks, as higher weights imply smaller HS distances.
Next, we propose a Q-CurL game to further examine the effect of Q-CurL. In this game, Alice has an ML model ${\mathcal{M}}({\bm{\theta}})$ to solve the main task ${\mathcal{T}}_{M}$, but she needs to solve all the auxiliary tasks ${\mathcal{T}}_{1},\ldots,{\mathcal{T}}_{M-1}$ first. We assume the data forgetting in task transfer, meaning that after solving task $A$, only the trained parameters ${\bm{\theta}}_{A}$ are transferred as the initial parameters for task $B$. We propose the following greedy algorithm to decide the curriculum order ${\mathcal{T}}_{i_{1}}\to{\mathcal{T}}_{i_{2}}\to\ldots\to{\mathcal{T}}_{i_{M}=M}$ before training. Starting ${\mathcal{T}}_{i_{M}}$, we find the auxiliary task ${\mathcal{T}}_{i_{M-1}}$ ($i_{M-1}\in\{1,2,\ldots,M-1\}$) with the highest curriculum weights $c_{i_{M},i_{M-1}}$. Similarity, to solve ${\mathcal{T}}_{i_{M-1}}$, we find the corresponding auxiliary task ${\mathcal{T}}_{i_{M-2}}$ in the remaining tasks with the highest $c_{i_{M-1},i_{M-2}}$, and so on. Here, curriculum weights $c_{i_{k},i_{k-1}}$ are calculated similarly to Eq. (5).
Figure 2: (a) The curriculum weight (lower panel) and the Hilbert-Schmidt distance (upper panel) between the target unitary of the main task ${\mathcal{T}}_{M}$ and the target unitary of the auxiliary task ${\mathcal{T}}_{m}$. (b) The training loss and test loss for different training epochs and different numbers $N$ of training data in the Q-CurL game, considering both random and Q-CurL orders. The average and standard deviations are calculated over 100 trials.
Figure 2(b) depicts the training and test loss of the main task ${\mathcal{T}}_{M}$ (see Eq. (6)) for different training epochs and numbers of training data over 100 trials of parameters’ initialization. In each trial, $N$ Haar random states are used for training, and 20 Haar random states are used for testing. With a sufficient amount of training data ($N=20$), introducing Q-CurL can significantly improve the trainability (lower training loss) and generalization (lower test loss) when compared with random order in Q-CurL game. Even with a limited amount of training data ($N=10$), when overfitting occurs, Q-CurL still performs better than the random order.
Data-based Q-CurL.— We present a form of data-based Q-CurL that dynamically predicts the easiness of each sample at each training epoch, such that easy samples are emphasized with large weights during the early stages of training and conversely. Remarkably, it does not involve pre-training or additional training data, thereby avoiding any increase in quantum resource requirements.
Apart from improving generalization, data-based Q-CurL offers resistance to noise. This feature is particularly valuable in QML, where clean annotated data are often costly while noisy data are abundant. Existing QML models can accurately fit corrupted labels in the training data but often fail on test data [25]. We demonstrate that data-based Q-CurL enhances robustness by dynamically weighting the difficulty of fitting corrupted labels.
Figure 3: The test loss and accuracy of the trained QCNN (with and without using the data-based Q-CurL) in the quantum phase recognition task with 8 qubits under varying noise levels in corrupted labels. Here, the average and the best performance over 50 trials are plotted.
Inspired by the confidence-aware techniques in classical ML [19, 20, 21], the idea is to modify the empirical risk as
$\hat{R}(h,{\bm{w}})=\dfrac{1}{N}\sum_{i=1}^{N}\left((\ell_{i}-\eta)e^{w_{i}}+\gamma w^{2}_{i}\right).$ (7)
Here, ${\bm{w}}=(w_{1},\ldots,w_{N})$, $\ell_{i}=\ell(h({\bm{x}}_{i}),{\bm{y}}_{i})$, and $w^{2}_{i}$ is the regularization term controlled by the hyper-parameter $\gamma>0$. The threshold $\eta$ distinguishes easy and hard samples with $e^{w_{i}}$ emphasizing the loss $l_{i}\ll\eta$ (easy sample) and neglecting the loss $l_{i}\gg\eta$ (hard samples, such as data with corrupted labels) [^1] In Supplementary Material, we have also discussed an interesting scenario where the modified loss in Eq. (7) can be used to emphasize complex quantum data during training, potentially reducing generation errors in quantum phase detection tasks under specific conditions. This aligns with the numerical results reported in Ref. [26], which appeared on arXiv after our paper.. The optimization is reduced to $\textup{min}_{{\bm{\theta}}}\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$, where ${\bm{\theta}}$ is the parameter of the hypothesis $h$. Here, $\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$ is decomposed at each loss $\ell_{i}$ and solved without quantum resources as $w_{i}=\textup{argmin}_{w}(l_{i}-\eta)e^{w}+\gamma w^{2}$. To control the difficulty of the samples, in each training epoch, we set $\eta$ as the average value of all $\ell_{i}$ obtained from the previous epoch. Therefore, $\eta$ adjusts dynamically in the early training stages but stabilizes near convergence.
We apply the data-based Q-CurL to the quantum phase recognition task investigated in Ref. [10] to demonstrate that it can improve the generalization of the learning model. Here, we consider a one-dimensional cluster Ising model with open boundary conditions, whose Hamiltonian with $Q$ qubits is given by $H=-\sum_{j=1}^{Q-2}\sigma^{z}_{j}\sigma^{x}_{j+1}\sigma^{z}_{j+2}-h_{1}\sum_{j=1}^{Q}\sigma^{x}_{j}-h_{2}\sum_{j=1}^{Q-1}\sigma^{x}_{j}\sigma^{x}_{j+1}.$ Depending on the coupling constants $(h_{1},h_{2})$, the ground state wave function of this Hamiltonian can exhibit multiple states of matter, such as the symmetry-protected topological phase, the paramagnetic state, and the anti-ferromagnetic state. We employ the quantum convolutional neural network (QCNN) model [10] with binary cross-entropy loss for training. Without Q-CurL, we use the conventional loss $\hat{R}(h)=(1/N)\sum_{i=1}^{N}\ell_{i}$ for the training and test phase. In data-based Q-CurL, we train the QCNN with the loss $\hat{R}(h,{\bm{w}})$ while using $\hat{R}(h)$ to evaluate the generalization on the test data set. We use 40 and 400 ground state wave functions for the training and test phases, respectively (see [23] for details).
We consider a scenario involving corrupted labels to evaluate the effectiveness of data-based Q-CurL in handling data difficulty during training. With a noise level probability $p$ ($0\leq p\leq 1$), the true label $y_{i}\in\{0,1\}$ of training state |$\ket{\psi_{i}}$ is flipped to the label $1-y_{i}$ with probability $p$, while it remains unchanged with probability $1-p$. Figure 3 illustrates the performance of a trained QCNN on test data across various noise levels. There is a minimal difference at low noise levels, but as noise increases, conventional training fails to generalize effectively. Introducing data-based Q-CurL in training (red lines) reduces test loss and improves test accuracy compared to the conventional method (blue lines). As further presented in  [23], Q-CurL enhances phase separation in the phase diagram, offering more reliable insights into the use of QML for understanding physical systems.
Discussion.— The proposed Q-CurL framework can enhance training convergence and generalization in QML with quantum data. Future research should investigate whether Q-CurL can be designed to improve trainability in QML, particularly by avoiding the barren plateau problem. For instance, curriculum design is not limited to tasks and data but can also involve the progressive design of the loss function. Even when the loss function of the target task, designed for infeasibility in classical simulation to achieve quantum advantage [27, 28], is prone to the barren plateau problem, a well-designed sequence of classically simulable loss functions can be beneficial. Optimizing these functions in a well-structured curriculum before optimizing the main function may significantly improve the trainability and performance of the target task.
Acknowledgements.
The authors acknowledge Koki Chinzei and Yuichi Kamata for their fruitful discussions. Special thanks are extended to Koki Chinzei for his valuable comments on the variations of the Q-CurL game, as detailed in the Supplementary Materials.